Determine the net yield of ATP from 10 glucose molecules released from muscle glycogen during anaerobic conditions and compare this to the net yield of ATP from 10 glucose molecules obtained from circulating blood.
Answer #1

Net ATP from 1 glucose molecule from circulating blood stream

Is 30 ATP

so from 10 glucose molecule total =30*10=300 ATP prosduced

Glucose release from glycogen produces =31 ATP

Because in glycogenolysis glycogen converted into glucose 1 phosphate and it is converted into Glucose 6 phosphate and then enter into glycolysis from here,,,, there is no ATP required yo convert glycogen into glucose 1 p or glucose 1 phosphate to glucose 6 phosphate (in glycolysis when glucose cinvert into glucose 6 phosphate ATP is required)

So there is 1 less consumption of ATP

So, here 31 ATP from 1 glucose from glycogen

From 10 glucose=31*10

=310 ATP